The enthalpy of fusion of ice is 6.01 kJ mol-1, and ice shrinks by 1.70 cm3 per mole, when it melts at 1 atm pressure. What would be the melting point of ice at 1cde atm (in my case, 1354 atm)? Assume 1 atm = 101325 Pa.


1
Expert's answer
2020-11-09T13:52:53-0500

Clausius-Clapeyron equation is:

dP/dT = "\Delta"H/T"\Delta"V

"\Delta"H = 6010 J/mole

"\Delta"V = -1.70*10-6 m3/mole

Then dT/T = dP*"\Delta"V/"\Delta"H

ln(T2/T1) = (P2-P1)*"\Delta"V/"\Delta"H

normal melting point is T1 = 273.15 K, P1 = 1 atm = 101325 Pa

At raised pressure P2 = 1354 atm the melting temperature T2 is:

T2 = T1*exp{(P2-P1)*"\Delta"V/"\Delta"H} = 273.15*exp{-1353*101325*1.70*10-6/6010} = 262.76 K

Melting point will be 262.76 K